OpenStudy (lovelycharm):

elimination -x + 5y = 133x + 5y = 27

OpenStudy (anonymous):

- x + 5y = 133 x + 5y = 27 ---------------- 10y = 160 y = 160/10 y = 16 now sub 16 in for y in either of the original equations... x + 5y = 27 x + 5(16) = 27 x + 80 = 27 x = 27 - 80 x = - 53 check... -x + 5y = 133 - (-53) + 5(16) = 133 53 + 80 = 133 133 = 133 (correct) x = - 53, y = 16 ---->(53,16) Do you have any questions about the problem I just did ?

OpenStudy (lovelycharm):

not yet

OpenStudy (lovelycharm):

thank u very much @kelliegirl33

OpenStudy (anonymous):

not a problem...glad to help :)